OpenStudy (anonymous):

x^2+8x+4=0 solve the equation (choose the best method)

OpenStudy (e.mccormick):

So what did you try?

OpenStudy (anonymous):

i can do the equation i just wanna know what method and in my book is that but i left in school :(

OpenStudy (e.mccormick):

When I look at \(x^2+8x+4=0\), the first thing I see is that they factors of 4 are either 1 and 4 or 2 and 2. There is no way those add up to 8. That means quadratic equation. Know that one?

OpenStudy (anonymous):

yep thanks :)

OpenStudy (anonymous):

Quadratic formula is the same as quadratic equation right

OpenStudy (e.mccormick):

\(ax^2+bx+c\) The relationship between \(a\cdot c\) and what adds up to b is what gives it away on these. When a is 1 it is not too hard.

OpenStudy (e.mccormick):

Yah.

OpenStudy (anonymous):

ok thanks

OpenStudy (e.mccormick):

This cute little thing:\[\frac{-b\pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

yep that's what i have

OpenStudy (e.mccormick):

OK, then that is what you need here.

OpenStudy (anonymous):

i'm gonna do it and i'm gonna give you my answer and u can check if its right

OpenStudy (e.mccormick):

OK.

OpenStudy (anonymous):

i got -0.26 and -7.74

OpenStudy (e.mccormick):

Do they want aprox. or exact values? I did exact values.

OpenStudy (anonymous):

sorry to bother you but what method is this one c^2-100=0

OpenStudy (e.mccormick):

I think you got your root wrong n that last one. In approx. I get -.54 -7.46

OpenStudy (anonymous):

exact

OpenStudy (anonymous):

ok thanks

OpenStudy (e.mccormick):

Exact means they will want it with the square root sign.

OpenStudy (anonymous):

ok :) so sorry but can you give me the method of c^2-100=0

OpenStudy (e.mccormick):

\[\frac{-8\pm \sqrt{8^2-4(1)(4)}}{2(1)} \implies \\ \frac{-8\pm \sqrt{64-16}}{2} \implies \\ \frac{-8\pm \sqrt{48}}{2} \implies \\ \frac{-8\pm 4\sqrt{3}}{2} \implies \\ -4\pm 2\sqrt{3}\] As for the second one, look at 100. What is another way of saying 100 that uses an exponent? If you know that, you should know which method.

OpenStudy (e.mccormick):

Recognize this form?\[c^2-100=0 \implies c^2-10^2=0\]

OpenStudy (anonymous):

no i don't remeber can u give me the name of the method tho?

OpenStudy (e.mccormick):

Difference of perfect squares

OpenStudy (e.mccormick):

\(a^2-b^2=(a+b)(a-b)\)

OpenStudy (anonymous):

oh like that i used another one

OpenStudy (anonymous):

c-100

OpenStudy (anonymous):

(c +100) (c-100)

OpenStudy (e.mccormick):

You have to solve both the + and the - side, and it is the non-squared one you want. So not 100, 10.

OpenStudy (e.mccormick):

Because \((10)(-10)=-100\)

OpenStudy (anonymous):

so c-100 and c 100

OpenStudy (anonymous):

i mean 10

OpenStudy (e.mccormick):

Yes, \(c=\pm 10\)

OpenStudy (e.mccormick):

Or in set notation, \(c=\{-10,10\}\)

OpenStudy (anonymous):

i bother way to much but if you are free can u say me the the method of 2x^2-54=0

OpenStudy (e.mccormick):

Well, 2 is not exactly a square... hmmm... neither is 54.... If you found some multiple in there that added up to 0, it might work... 54=9*6=3*3*3*2, so you might be able to just factor out the 2 and see if there is some other common factor in there...

OpenStudy (e.mccormick):

I don't see that as one I would just say, "This method will work."

OpenStudy (anonymous):

\[2x^2=54 \]

OpenStudy (anonymous):

\[x^2=54\div2\]

OpenStudy (anonymous):

\[\sqrt{49}\]

OpenStudy (anonymous):

x\[x =\pm 7\]

OpenStudy (anonymous):

the teacher taught us that one but idk if its correct

OpenStudy (anonymous):

bye and thank :*

OpenStudy (e.mccormick):

Well, the easist might be to factor the 2 and then just do difference of sqares, but then you are using a somewhat odd number for the b part. Still, it would work.

OpenStudy (e.mccormick):

54/2=27

OpenStudy (e.mccormick):

\[2x^2-54=0\implies \\ 2(x^2-27)=0\implies \\ (x+\sqrt{27})(x-\sqrt{27})=0\implies \\ x=\pm\sqrt{27} \]

OpenStudy (e.mccormick):

Or with your teacher's way: \[2x^2-54=0 \implies \\ 2x^2=54 \implies \\ x^2=27 \implies \\ x=\sqrt{27}\implies \\ x=\pm 3\sqrt{3}\] Which is the same thing.